# Prigogine’s temporalization of physics: two agnostic attitudes of physicists

### v2

@@ 3 Concept of time in Physics II: the time with or without process
When considering the description of change as a central axis, it becomes apparent that distinct classifications of change exist. Abstracting the notion of change reveals three fundamental phases: the initial state, the intermediate state, and the final state (See Fig. 1). For instance, in the phase transition of ice melting at room temperature, we can clearly identify the initial state as ice, the intermediate state as the melting process, and the final state as liquid water. Thermodynamics provides a rigorous theoretical framework for predicting such state transitions, but the specific elements included in the analysis lead to distinct theoretical formulations. Since the introduction of the second law of thermodynamics by Clausius, equilibrium thermodynamics has been established as a theory that exclusively deals with the relationship between initial and final states, while remaining agnostic about the intermediate state. This means that equilibrium thermodynamics can rigorously answer the question, ”What happens after ice is placed at room temperature?” but fails to address questions concerning ”How does the ice melt?” or ”When does the melting occur?”. Theories that account for intermediate states require an entirely different framework, namely kinetics [21, 22]  [^2] Soviet physicists tends to call the non-equilibrium theory of statistical physics a term, Physical Kinetics [21, 22]. and non-equilibrium thermodynamics [23, 19]. Crucially, these theories incorporate both entropy production and the concept of time to achieve a more comprehensive description of dynamical processes.
@@
Notably, the two frameworks of equilibrium and non-equilibrium thermodynamics did not emerge simultaneously. The equilibrium thermodynamics was established in the mid-19th century [24], whereas the non-equilibrium thermodynamics only emerged after 1931  [^3] The statement that non-equilibrium thermodynamics began in 1931 refers to the fact that the explore for universally applicable laws governing all non-equilibrium states, although they were limited to the linear regime, began after the discovery of Onsager’s reciprocal relations in 1931. Directly, it means that entropy production was successfully directly estimated. However, the study of non-equilibrium phenomena themselves dates back much earlier.. Prior to this, classical dynamics and equilibrium thermodynamics constituted the foundational theories, and the later development of non-equilibrium thermodynamics represented an effort to unify these two domains.
@@
Examining non-equilibrium thermodynamics in the nonlinear regime presents an intriguing perspective. Non-equilibrium thermodynamics can be further categorized into two domains: linear non-equilibrium thermodynamics and nonlinear non-equilibrium thermodynamics (See Fig. 2). The former applies to regions near equilibrium, where the assumption of linearity between the forces and flows holds [19]. Although it deals with non-equilibrium systems, linear non-equilibrium thermodynamics operates under the principle of minimum entropy production, ensuring that the system evolves toward a state that minimizes entropy production, given the imposed boundary conditions. This implies the existence of a well-defined final state. However, in the nonlinear regime, such a final state may cease to exist under certain conditions. This phenomenon was first identified by Prigogine and colleagues in the Belousov–Zhabotinsky (BZ) reaction and was later demonstrated in a model known as the Brusselator. In the Brusselator model, an autocatalytic reaction pathway plays a crucial role. In conventional chemical reactions, reactants and products are molecularly distinct, as in the case of $A+B\rightarrow C+D$. However, in an autocatalytic reaction, the reactants and products include the same molecular species, as exemplified by $2X+Y\rightarrow 3X$. Mathematically, this results in a nonlinear term of second order in the reaction rate equation, forming a feedback loop where the reaction’s outcome (final state) influences its cause (initial state). This distinguishes nonlinear systems from linear ones, where cause and effect are clearly separable. In nonlinear regimes, however, these elements become entangled. Such reaction pathways create conditions in which instabilities are amplified, leading to emergent spatial and spatiotemporal patterns under appropriate conditions [^4] In the linear regime, the linear relation of the phenomenological force and the flow, the minimum entropy production and the reciprocal relations are generally satisfied to get Lyapnov function while, in the nonlinear regime, these linearities are broken, then the Lyapunov function is not satisfied. See Ch.18, 19 of Ref. [19]. When spatial diffusion processes are involved, periodic spatial oscillations and self-organized pattern formation can occur, giving rise to structured spatial or spatiotemporal dynamics.
@@
While these classifications do not strictly separate different scientific disciplines, they allow for a characterization of the concept of time in physics. This characterization is based on whether the three essential properties—difference, process, and monotonicity—are present or absent. In other words, the concept of time in physics can be understood as an extended formulation of the aspects of time that we experience in daily life. A significant trend emerges from this analysis. In the early 20th century, the predominant physical theories were dynamics and equilibrium thermodynamics. Historically, non-equilibrium thermodynamics emerged later than equilibrium thermodynamics, and while the foundational theories of equilibrium thermodynamics have been well established, the fundamental theories of non-equilibrium thermodynamics remain in development. In particular, nonlinear non-equilibrium thermodynamics has yet to establish universal laws comparable to those in equilibrium thermodynamics and linear non-equilibrium thermodynamics. This development on the asymmetry seems to reflect the problems and fields in which physics demonstrates particular strengths and its limitations.
@@ 4 The philosophy of dynamics and thermodynamics
In contrast, the physics of being, which is fundamentally based on reversible processes—namely, dynamics—does not regard dissipation as a central problem. Dynamics was originally developed from kinematics, which aimed to describe motion without considering its causes, focusing instead on its mathematical structure. It was Newton who introduced causal forces, such as universal gravitation, into the study of motion. Following Newton, Euler, Lagrange, and other continental mathematicians refined and systematized the field [30]. The origins of dynamics are therefore deeply rooted in mathematics, and its fundamental equations naturally emerge from principles such as time homogeneity and spatial homogeneity. It focuses on the conservation of energy, treating work as a well-defined, quantifiable entity. In the fundamental equations of dynamics, no unknown quantity corresponding to dissipation exists. Consequently, these equations predict motion symmetrically in both the past and the future. Thus, within the fundamental laws of dynamics, there was never a logical necessity to incorporate dissipation into its foundational equations.
@@
In contrast, thermodynamics treats dissipation as a crucial concept that is explicitly incorporated into its fundamental equations, most notably through the second law of thermodynamics. The first law represents the principle of energy conservation, while the second law establishes the directionality of natural processes and asserts the existence of irreversibility. In this sense, thermodynamics can be regarded as the science of dissipation. Thus, dissipation occupies fundamentally different positions in the two major fields of physics. It is important to note that dissipation is indeed considered in dynamics. However, the crucial distinction is that, in dynamics, dissipation is treated as a convenient approximation-a secondary modification added to idealized equations to align with empirical reality. At its core, it is viewed as a redundant or extraneous concept rather than a fundamental principle.
@@ 5 Naive agnostic attitude and essential agnostic attitude
From Huxley’s formulation, one can discern an agnostic stance as a defining feature of the scientific attitude—namely, an epistemic modesty that refrains from making definitive claims about matters not grounded in mathematical reasoning or empirical observation. A similar agnostic stance can be found in transcendental idealism. To withhold judgment on matters not secured by mathematical or empirical justification may thus be regarded as an expression of agnosticism. In this respect, physics—being grounded in theories that rely on experimental observation—necessarily entails intrinsic limits on the scope within which its claims can be considered valid. However, as we have seen in the previous section, the way in which these limits are positioned differs between dynamics and thermodynamics.
@@ 6 Spatialization of physics
To clarify this point, Bergson introduces the concept of durée—translated as duration—as a more faithful expression of the nature of time. Duration emphasizes time not as a series of discrete, measurable intervals, but as a flowing continuity and as a process unfolding in real time. In this view, the essence of time lies in its incompleteness and openness, which is something in the process of happening. Time as duration, in Bergson’s philosophy, is not a sequence of positions in space but a real and continuous becoming, an unfinished process that constitutes the very nature of time itself.
@@ 7 Temporalization of physics
However, this does not mean that the classification of pitch as either a solid or a liquid is purely subjective. The determining factor ultimately involves the relaxation time of the material itself, meaning that the physical properties of the substance must also be taken into account. What the pitch drop experiment fundamentally reveals is that the properties of a material are determined by the scale at which the subject (observer) and the object (material) interact. In this sense, the observer is inherently complicit in defining the attributes of pitch. This complicit  [^11] The term complicit here is used to highlight the modest contribution of the subject in determining the properties of materials. relation between observer and material is the fundamental insight offered by the pitch drop experiment.
For instance, glass has a relaxation time of 19,000 years [69]. Following the same logic, glass would exhibit fluid-like behavior on a 10,000-year time scale. However, for humans, glass is perceived as a solid, because in our daily experience, glass does not flow. The reason for this is that the human interaction time with glass typically ranges from a few hours to at most a lifetime of several decades. Within this human time scale, glass has not fully manifested its potential behaviors, yet what is meaningful for humans is the glass as it appears within the constraints of human finiteness—as something in the midst of a process that is imperceptible to us. In this sense, humans play a role in determining the perceived properties of glass.  [^12] Glass has long been a subject of research within the physics community and remains a topic of ongoing debate. The present formulation is grounded in the concept of the Deborah number and the framework of intermediate asymptotics..
@@
Of course, Bergson’s philosophy and Prigogine’s work do not completely overlap. Prigogine engaged with the concept of time within the realm of physics, whereas Bergson dealt with a broader critique of science, metaphysics, and philosophy, focusing on the fundamental errors that human intellect inevitably makes when attempting to conceptualize time. In this sense, Bergson and Prigogine operate on different levels. Nevertheless, their projects share a common goal. Just as Bergson sought to restore the true nature of time, challenging both science and philosophy for reducing it to a static, discontinuous entity, Prigogine sought to restore time within physics itself. In the physics of being, dissipation and time were regarded as excessive, incomplete, and fundamentally negative concepts. Prigogine’s work aimed to reposition them in their rightful place, restoring their intrinsic significance within physics.

### v3

@@ 3 Concept of time in Physics II: the time with or without process
When considering the description of change as a central axis, it becomes apparent that distinct classifications of change exist. Abstracting the notion of change reveals three fundamental phases: the initial state, the intermediate state, and the final state (See Fig. 1). For instance, in the phase transition of ice melting at room temperature, we can clearly identify the initial state as ice, the intermediate state as the melting process, and the final state as liquid water. Thermodynamics provides a rigorous theoretical framework for predicting such state transitions, but the specific elements included in the analysis lead to distinct theoretical formulations. Since the introduction of the second law of thermodynamics by Clausius, equilibrium thermodynamics has been established as a theory that exclusively deals with the relationship between initial and final states, while ignoring the intermediate state. This means that equilibrium thermodynamics can rigorously answer the question, ”What happens after ice is placed at room temperature?” but fails to address questions concerning ”How does the ice melt?” or ”When does the melting finish?”. Theories that account for intermediate states require an entirely different framework, namely kinetics [21, 22]  [^2] Soviet physicists tends to call the non-equilibrium theory of statistical physics a term, Physical Kinetics [21, 22]. and non-equilibrium thermodynamics [23, 19]. Crucially, these theories incorporate both entropy production and the concept of time to achieve a more comprehensive description of dynamical processes.
@@
Notably, the two frameworks of equilibrium and non-equilibrium thermodynamics did not emerge simultaneously. The equilibrium thermodynamics was established in the mid-19th century [24], whereas the non-equilibrium thermodynamics only emerged after 1931  [^3] The statement that non-equilibrium thermodynamics began in 1931 refers to the fact that the explore for universally applicable laws governing all non-equilibrium states, although they were limited to the linear regime, began after the discovery of Onsager’s reciprocal relations in 1931. However, the study of non-equilibrium phenomena themselves dates back much earlier.. Prior to this, classical dynamics and equilibrium thermodynamics constituted the foundational theories, and the later development of non-equilibrium thermodynamics represented an effort to unify these two domains.
@@
Examining non-equilibrium thermodynamics in the nonlinear regime presents an intriguing perspective. Non-equilibrium thermodynamics can be further categorized into two domains: linear non-equilibrium thermodynamics and nonlinear non-equilibrium thermodynamics (See Fig. 2). The former applies to regions near equilibrium, where the assumption of linearity between the forces and flows holds [19]. Although it deals with non-equilibrium systems, linear non-equilibrium thermodynamics operates under the principle of minimum entropy production, ensuring that the system evolves toward a state that minimizes entropy production, given the imposed boundary conditions. This implies the existence of a well-defined final state. However, in the nonlinear regime, such a final state may cease to exist under certain conditions. This phenomenon was first identified by Prigogine and colleagues in the Belousov–Zhabotinsky (BZ) reaction and was later demonstrated in a model known as the Brusselator. In the Brusselator model, an autocatalytic reaction pathway plays a crucial role. In conventional chemical reactions, reactants and products are molecularly distinct, as in the case of $A+B\rightarrow C+D$. However, in an autocatalytic reaction, the reactants and products include the same molecular species, as exemplified by $2X+Y\rightarrow 3X$. Mathematically, this results in a nonlinear term of second order in the reaction rate equation, forming a feedback loop where the reaction’s outcome (final state) influences its cause (initial state). This distinguishes nonlinear systems from linear ones, where cause and effect are clearly separable. In nonlinear regimes, however, these elements become entangled. Such reaction pathways create conditions in which instabilities are amplified, leading to emergent chemical oscillations under appropriate conditions [^4] In the linear regime, the linear relation of the phenomenological force and the flow, the minimum entropy production and the reciprocal relations are generally satisfied to get Lyapnov function while, in the nonlinear regime, these linearities are broken, then the Lyapunov function is not satisfied. See Ch.18, 19 of Ref. [19]. When spatial diffusion processes are involved, self-organized pattern formation can occur, giving rise to structured spatial or spatiotemporal dynamics, which is dissipative structure.
@@
While these classifications do not strictly separate different scientific disciplines, they allow for a major characterization of the concept of time in each fields. This characterization is based on whether the three essential properties—difference, process, and monotonicity—are present or absent. In other words, the concept of time in physics can be understood as an extended formulation of the aspects of time that we experience in daily life. A significant trend emerges from this analysis. In the early 20th century, the predominant physical theories were dynamics and equilibrium thermodynamics. Historically, non-equilibrium thermodynamics emerged later than equilibrium thermodynamics, and while the foundational theories of equilibrium thermodynamics have been well established, the fundamental theories of non-equilibrium thermodynamics remain in development. In particular, nonlinear non-equilibrium thermodynamics has yet to establish universal laws comparable to those in equilibrium thermodynamics and linear non-equilibrium thermodynamics. The asymmetry in the development of these fields arises from the characteristics of the problems intrinsic to each field and reflects how time is conceptualized.
@@ 4 The philosophy of dynamics and thermodynamics
In contrast, the physics of being, which is fundamentally based on reversible processes—namely, dynamics—does not regard dissipation as a central problem. Dynamics was originally developed from kinematics, which aimed to describe motion without considering its causes, focusing instead on its mathematical formulation. It was Newton who introduced causal forces, such as universal gravitation, into the study of motion. Following Newton, Euler, Lagrange, and other continental mathematicians refined and systematized the field [30]. The origins of dynamics are therefore deeply rooted in mathematics, and its fundamental equations naturally emerge from principles such as time homogeneity and spatial homogeneity. It focuses on the conservation of energy, treating work as a well-defined, quantifiable entity. In the fundamental equations of dynamics, no unknown quantity corresponding to dissipation exists. Consequently, these equations predict motion symmetrically in both the past and the future. Thus, within the fundamental laws of dynamics, there was never a logical necessity to incorporate dissipation into its foundational equations.
@@
In contrast, thermodynamics treats dissipation as a crucial concept that is explicitly incorporated into its fundamental equations, most notably through the second law of thermodynamics. The first law represents the principle of energy conservation, while the second law establishes the directionality of natural processes and asserts the existence of irreversibility. In this sense, thermodynamics can be regarded as the science of dissipation. Thus, dissipation occupies fundamentally different positions in the two major fields of physics. It is important to note that dissipation is indeed considered in dynamics. However, the crucial distinction is that, in dynamics, dissipation is treated as a convenient approximation—a secondary modification added to idealized equations to align with empirical reality. At its core, it is viewed as a redundant or extraneous concept rather than a fundamental principle.
@@ 5 Naive agnostic attitude and essential agnostic attitude
From Huxley’s formulation, one can discern an agnostic stance as a defining feature of the scientific attitude—namely, an epistemic modesty that refrains from making definitive claims about matters not grounded in mathematical reasoning or empirical observation. In this respect, physics—being grounded in theories that rely on experimental observation—necessarily entails intrinsic limits on the scope within which its claims can be considered valid. This suggests the existence of an intrinsic attitude or strategy toward its inherent limitations. As we have seen in the previous section, the way in which these limits are positioned differs between dynamics and thermodynamics.
@@ 6 Spatialization of physics
To clarify this point, Bergson introduces the concept of durée—translated as duration—as the nature of time. Duration emphasizes time not as a series of discrete, measurable intervals, but as a flowing continuity and as a process unfolding in real time. In this view, the essence of time lies in its incompleteness and openness, which is something in the process of happening. Time as duration, in Bergson’s philosophy, is not a sequence of positions in space but a real and continuous becoming, an unfinished process that constitutes the very nature of time itself.
@@ 7 Temporalization of physics
However, this does not mean that the classification of pitch as either a solid or a liquid is purely subjective. The determining factor ultimately involves the relaxation time of the material itself, meaning that the physical properties of the substance must also be taken into account. What the pitch drop experiment fundamentally reveals is that the properties of a material are determined by the scale at which the subject (observer) and the object (material) interact. In this sense, the observer is inevitably complicit in defining the attributes of pitch. This complicit  [^11] The term complicit here is used to highlight the inevitable and inseparable contribution of the subject in determining the properties of materials. relation between observer and material is the fundamental insight offered by the pitch drop experiment.
For instance, glass has a relaxation time of 19,000 years [69]. Following the same logic, glass would exhibit fluid-like behavior on a 10,000-year time scale. However, for humans, glass is perceived as a solid, because in our daily experience, glass does not flow. The reason for this is that the human interaction time with glass typically ranges from a few hours to at most a lifetime of several decades. Within this human time scale, glass has not fully manifested its potential behaviors, yet what is meaningful for humans is the glass as it appears within the constraints of human finiteness—as something in the midst of a process that is involving us. In this sense, humans play a role in determining the perceived properties of glass.  [^12] Glass has long been a subject of research within the physics community and remains a topic of ongoing debate. The present formulation is grounded in the concept of the Deborah number and the framework of intermediate asymptotics..
@@
Of course, Bergson’s philosophy and Prigogine’s work do not completely overlap. Prigogine engaged with the concept of time within the realm of physics, whereas Bergson dealt with a broader critique of science, metaphysics, and philosophy, focusing on the fundamental errors that human intellect inevitably makes when attempting to conceptualize time. In this sense, Bergson and Prigogine operate discuss at different levels. Nevertheless, their projects share a common goal. Just as Bergson sought to restore the true nature of time, challenging both science and philosophy for reducing it to a static, discontinuous entity, Prigogine sought to restore time within physics itself. In the physics of being, dissipation and time were regarded as excessive, incomplete, and fundamentally negative concepts. Prigogine’s work aimed to reposition them in their rightful place, restoring their intrinsic significance within physics.